OpenStudy (anonymous):

A 0.5 kg block is moving at constant speed on a horizontal surface. the friction force acting on the block is 20N, calculate uk

OpenStudy (anonymous):

*uk is the coefficient of friction

OpenStudy (anonymous):

would it just be .5 *20

OpenStudy (anonymous):

any idea?

OpenStudy (cwrw238):

oh right the opposing forces are equal as the speed is constant friction al force = mN where N = normal reaction of ground on the block = m* 0.5g = 20 m = 20 / (0.5*9.81)

OpenStudy (cwrw238):

hmmm - can't understand this - the uk must be in range 0 - 1. maybe the values in question are wrong

OpenStudy (anonymous):

so is your work right? whats its units?

OpenStudy (cwrw238):

i'm pretty sure the methods right - units???

OpenStudy (anonymous):

cause the answer comes out to be 4.08? so 4.08 Neutons or what

OpenStudy (cwrw238):

not sure - ill check

OpenStudy (anonymous):

ok?

OpenStudy (cwrw238):

no units as its a coefficient

OpenStudy (anonymous):

so that answer is correct?

OpenStudy (cwrw238):

the answer is correct given the values - but in practice a force of 20 newtons is not possible for a 0.5 kg mass.

OpenStudy (anonymous):

the question i typed is what it says on my paper...

OpenStudy (cwrw238):

frictional force = uk * normal reaction of ground on the block normal reaction = mass * g = 0.5 * 9.81

OpenStudy (cwrw238):

now 2 N would be a possible value and uk would then be 0.408

OpenStudy (anonymous):

can you explain why its .408

OpenStudy (cwrw238):

right frictional force = 2 = 0.5*9.81 * uk uk = 2 / (0.5*9.81) = 0.408

OpenStudy (anonymous):

can you give me a medal too man?

OpenStudy (cwrw238):

of course if the mass is 5 kg then 20 N would be the force - that would give the same answer